OpenStudy (anonymous):

For the function , make a table with integer values of x from 0 to 4. Then graph the function (you do not need to submit the graph). Does the graph of show exponential growth, exponential decay, or neither? Explain your thinking.

OpenStudy (anonymous):

\[y =1^{x}\]

OpenStudy (ddcamp):

Fill in the graph and plot the points:  If you want, clicking the pen on the top corner of the drawing will allow you to draw over what's already there.

OpenStudy (hunus):

What are the values of y in these cases?\[y=1^0\]\[y=1^1\]\[y=1^2\]\[y=1^3\]\[y=1^4\]

OpenStudy (anonymous):

Im trying to understand what integer values of x are.

OpenStudy (e.mccormick):

Integer, the numbers without decimals or fractions.

OpenStudy (ddcamp):

So you only test whole numbers (1,2,3,-1,-5,0, to name a few)

OpenStudy (e.mccormick):

Huh? -1 and -5 are not whole nunbers.

OpenStudy (anonymous):

but x = 1 isn't it ture that 1^x is always 1

OpenStudy (hunus):

Yes

OpenStudy (e.mccormick):

Yes GertheDL, they are all 1. What does that look like? Exponential growth, exponential decay, or neither?

OpenStudy (anonymous):

so there is no growth or decay?

OpenStudy (hunus):

Does y ever equal anything besides 1?

OpenStudy (anonymous):

nope

OpenStudy (e.mccormick):

Yep GertheDL! It is neither. -7, 259, 42 and things like that are all integers. Where as \(\frac{1}{3}\) is a rational number (from ratio). In contrast, \(\pi\), \(e\), and \(\sqrt{2}\) are irrational as in they can not be represented by a ratio. There are also complex, whole, counting, and a few others. Real numbers is an important category. Whole numbers are the integers \(\ge 0\).

OpenStudy (hunus):

So it doesn't grow or decay :)

OpenStudy (e.mccormick):

When you do the explain your thinking part, what you said about \(1^x=1 \forall x \in \mathbb{Z} \ge 0\) will be good.